# The finite groups with three automorphism orbits

### v1

@@ 2.3. Extraspecial $q$-groups
Thus, $w^{g}\in W$ for any $w\in W$ and $g\in G$. Hence, $G$ is a subgroup of ${\mathrm{Aut}}(N)^{V}$. By Lemmas 2.8 (i) and 2.11, the preimage $A$ of $G$ in ${\mathrm{Aut}}(N)$ is isomorphic to $p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$, and $A^{V}=G\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$.
@@ 3. Exterior Square of Transitive Linear Groups
Let $p$ be an odd prime, $V$ be a vector space over ${\mathbb{F}}_{p}$. To classify special $3$-orbit $p$-group, Lemma 2.8 (iii) and (iv) lead us to focus on the pairs $(G,W)$ such that $W<\Lambda_{{\mathbb{F}}_{p}}^{2}(V)$ is invariant under $G\leqslant{\mathrm{GL}}(V)$ and $G$ acts transitive on non-zero vectors of both $V$ and $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)/W$.
We remark that a linear group $G\leqslant{\mathrm{GL}}(V)$ acting transitively on non-zero vectors is called a transitive linear group. The classification of transitive linear groups is a part of well-known classification of $2$-transitive groups. The following theorem can be concluded from [6, Table 7.3].
@@
Let $G\leqslant{\mathrm{GL}}(n,p)$ be a transitive linear group with odd prime $p$. Then one of the following statements hold.
@@
Let $U$ be an ${\mathrm{E}}G$-module corresponding to the homomorphism $\rho:G\rightarrow{\mathrm{GL}}(V)$. For $\phi\in{\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$, denote by $U^{\phi}$ the ${\mathrm{E}}G$-module such that $u^{g}:=(u^{\rho(g)})^{\phi}$ for $u\in U$ and $g\in G$. For an ${\mathrm{F}}G$-module $V$, $V^{\mathrm{E}}:=V\otimes_{\mathrm{F}}{\mathrm{E}}$ is naturally an ${\mathrm{E}}G$-module. The following lemma on decompositions of exterior squares is [22, Lemma 2.2].
@@
Suppose that $V$ is an ${\mathrm{F}}G$-module and ${\mathrm{E}}$ is an extension field of ${\mathrm{F}}$. If $U$ is an ${\mathrm{E}}G$-module such that $U\cong V$ as ${\mathrm{F}}G$-modules, then the following statements hold.
@@
    $V^{\mathrm{E}}\cong\bigoplus_{i=1}^{f}U^{\phi^{i}}$ where $\phi$ is a generator of ${\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$ and $f=[{\mathrm{E}}:{\mathrm{F}}]$;
@@
    $\Lambda^{2}_{\mathrm{F}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$ where
    1. (1)
      $A\cong\Lambda_{\mathrm{E}}^{2}(U)$ as ${\mathrm{F}}G$-modules;
    2. (2)
    3. (3)
The following lemma is deduced by [22, Lemma 2.6] based on Steinberg’s Twisted Tensor Product Theorem [28] (also see [18, Theorem 5.4.1]).
Lemma 3.3.
Suppose that ${\mathrm{SL}}(m,q)\leqslant G\leqslant{\mathrm{GL}}(m,q)$ or ${\mathrm{Sp}}(m,q)\leqslant G\leqslant{\mathrm{CSp}}(2m,q)$ with $q=p^{f}$. Let $V$ and $U$ be the natural ${\mathbb{F}}_{p}G$-module and the natural ${\mathbb{F}}_{q}G$-module, respectively. Then $U\cong V$ as ${\mathbb{F}}_{p}G$-modules, and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$ where
  1. (i)
    $A\cong\Lambda_{{\mathbb{F}}_{q}}^{2}(U)$ as ${\mathbb{F}}_{p}G$-modules;
  2. (ii)
    for $j<f/2$, $B_{j}\cong U\otimes_{{\mathbb{F}}_{q}}U^{\phi^{j}}$ as irreducible ${\mathbb{F}}_{p}G$-module, and $B_{j}$’s are pairwise non-isomorphic;
@@
    when $f$ is even, $2\cdot B_{f/2}\cong U\otimes_{{\mathbb{F}}_{q}}U^{\phi^{f/2}}$ as ${\mathbb{F}}_{p}G$-modules, and $B_{f/2}$ is an irreducible ${\mathbb{F}}_{p}G$-module.
The exterior squares of natural modules for classical groups are well-known. We state the following useful results for ${\mathrm{SL}}(m,q)$ below.
Lemma 3.4 ([23, Theorem 2.2]).
Let $U={\mathbb{F}}_{q}^{m}$ with $q=p^{f}$, and let $G={\mathrm{SL}}(m,q)$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an irreducible ${\mathbb{F}}_{q}G$-module, and it is also an irreducible ${\mathbb{F}}_{p}G$-module.
In particular, $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is isomorphic to the dual of $U$ when $m=3$.
With the decomposition given in Lemma 3.3, we obtain the following lemma.
Lemma 3.5.
Suppose that $G\leqslant{\mathrm{GL}}(n,p)$ has a normal subgroup $G_{0}\cong{\mathrm{SL}}(m,p^{f})$ with $p$ a prime, $n=mf$ and $m\geqslant 3$. Let $V$ be the natural ${\mathbb{F}}_{p}G$-module. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ for ${\mathbb{F}}_{p}G$-submodules $A$ and $B$ such that
  1. (i)
    $A\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(U)$ as ${\mathbb{F}}_{p}G_{0}$-modules where $U$ is the natural ${\mathbb{F}}_{q}G_{0}$-module; and
  2. (ii)
    $G$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ if and only if $W=B$ and $m=3$.
@@
By Lemma 3.3, we can decompose $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ to be a direct sum of ${\mathbb{F}}_{p}G_{0}$-submodules:
@@
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}.$
Let $B=\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ and part (i) holds by Lemma 3.3. Now we prove that both $A$ and $B$ are ${\mathbb{F}}_{p}G$-submodules. For $g\in G$, $\langle B^{g},B\rangle/B$ is an ${\mathbb{F}}_{p}G_{0}$-submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/B\cong A$. Since $A$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module by Lemma 3.4, either $B^{g}=B$ or $A$ is isomorphic to one of $B_{j}$’s for $1\leqslant j\leqslant\lfloor\frac{f}{2}\rfloor$ as each $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module by Lemma 3.3. We can obtain that $\dim A=\frac{m(m-1)f}{2}$ and $\dim B_{j}=\left\{\begin{aligned} &m^{2}f,j\neq f/2;\\ &m^{2}f/2,j=f/2.\end{aligned}\right.$. Hence, $\dim A\neq\dim B_{j}$ for each $j$, and thus $B^{g}=B$ for any $g\in G$. This yields that $B$ is an ${\mathbb{F}}_{p}G$-submodule. On the other hand, if $A^{g}\neq A$, then $A$ is isomorphic to one of $B_{j}$’s, which is impossible. Thus, $A$ is also an ${\mathbb{F}}_{p}G$-submodule.
If $W=B$ and $m=3$, then $A$ is isomorphic to the dual of $V$ as ${\mathbb{F}}_{p}G_{0}$-modules by Lemma 3.4. Hence, $G_{0}$ acts transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/B\cong A$, and so does $G$. Conversely, suppose that $G$ acts transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. Let $K$ be the kernel of $G$ acting on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. Note that $A$ and $B_{j}$’s are non-trivial irreducible ${\mathbb{F}}_{p}G_{0}$-submodules. Thus, $K\cap{\mathbf{Z}}(G_{0})\leqslant{\mathbf{Z}}(G_{0})$ and $G/K$ is non-solvable. By Theorem 3.1, we obtain that $K\cap{\mathbf{Z}}(G_{0})=1$, $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W=\dim V$ and $G_{0}$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. If $B\not\subseteq W$, then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\cong B_{j}$ as ${\mathbb{F}}_{p}G_{0}$-modules for some $21\leqslant j\leqslant f/$ 2. Since $m\geqslant 3$, we obtain that $\dim B_{j}\geqslant mf=\dim V$. Thus, $B\subseteq W$, and then $B=W$ and $A\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. Since $\dim A=\dim V$, we have $m=3$ as desired. ∎
Results of exterior square of the natural modules of symplectic groups can be found in [7, 19, 25], and we state them below.
Lemma 3.6.
Let $X={\mathrm{CSp}}(2m,q)$ and $G={\mathrm{Sp}}(2m,q)\lhd X$ acing naturally on symplectic space $U={\mathbb{F}}_{q}^{2m}$ with $q=p^{f}$ for odd prime $p$. Then $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ has a maximal ${\mathbb{F}}_{q}G$-submodule $T$ such that
  1. (i)
    $\dim_{{\mathbb{F}}_{q}}T=\dim_{{\mathbb{F}}_{q}}\Lambda^{2}_{{\mathbb{F}}_{q}}(U)-1$;
  2. (ii)
    $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)/T$ is a faithful ${\mathbb{F}}_{p}(X/G)$-module.
Moerover, $T$ is indecomposable and one the following statements holds.
  1. (i)
    When $p\nmid m$, $T$ is an irreducible ${\mathbb{F}}_{p}G$-module.
  2. (ii)
    When $p\mid m$, $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$ is an indecomposable ${\mathbb{F}}_{q}G$-module, $T$ has a unique maximal ${\mathbb{F}}_{q}G$-submodule $T_{0}$ which is trivial, and $T/T_{0}$ is an irreducible ${\mathbb{F}}_{p}G$-module.
Hence, we can deal with $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)$ for the symplectic case.
Lemma 3.7.
Suppose that $G\leqslant{\mathrm{GL}}(n,p)$ has a normal subgroup $G_{0}\cong{\mathrm{Sp}}(2m,q)$ with $p$ an odd prime, $q=p^{f}$, $n=2mf$ and $(2m,q)\neq(2,3)$. Let $V$ be the natural ${\mathbb{F}}_{p}G$-module. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ for ${\mathbb{F}}_{p}G$-submodules $A$ and $B$ and the following statements hold.
  1. (i)
    $A\cong\Lambda^{2}_{{\mathbb{F}}_{p}}(U)$ as ${\mathbb{F}}_{p}G_{0}$-modules where $U$ is the natural ${\mathbb{F}}_{q}G_{0}$-module.
  2. (ii)
    $A$ has an ${\mathbb{F}}_{p}G$-submodule $T$ such that $\dim A/T=f$ and $G_{0}$ is the kernel of $G$ acting on $A/T$.
  3. (iii)
    $S:=T\oplus B$ is an ${\mathbb{F}}_{p}G$-submodule and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$ if and only if $S\subseteq W$.
  4. (iv)
    $G$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ if and only if $S\subseteq W$ and $G/G_{0}\lesssim{\Gamma\mathrm{L}}(1,q)$ acting transitively on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$.
Proof.
By Lemma 3.3, we can decompose $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ to be a direct sum of ${\mathbb{F}}_{p}G_{0}$-submodules:
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}.$
Let $B=\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}$. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B$ and part (i) holds by Lemma 3.3. Now we show that $A$ and $B$ are ${\mathbb{F}}_{p}G$-submodules. Note that if $B^{g}\neq B$ for some $g\in G$, then $A$ has an irreducible ${\mathbb{F}}_{p}G_{0}$-submodule whose dimension equals to $\dim B_{j}$ for some $1\leqslant j\leqslant\lfloor\frac{f}{2}\rfloor$. By Lemma 3.6, irreducible ${\mathbb{F}}_{p}G_{0}$-submodules of $A$ has dimension equal to one of $\left(\binom{2m}{2}-1\right)f$ and $\left(\binom{2m}{2}-2\right)f$, and none of them equal to $\dim B_{j}$ for any $j$. Hence, $B^{g}=B$ for any $g\in G$ and $B$ is an ${\mathbb{F}}_{p}G$-module. Similarly, we obtain that $A$ is also an ${\mathbb{F}}_{p}G$-submodule. By Lemma 3.6, $A$ has a maximal ${\mathbb{F}}_{q}G_{0}$-submodule $T$ such that $\dim_{{\mathbb{F}}_{p}}A/T=f$. If $T^{g}\neq T$, then $\langle T,T^{g}\rangle/T$ is a non-trivial $F_{p}G_{0}$=submodule of $A/T$, which is a contradiction. Thus, $T$ is an ${\mathbb{F}}_{p}G$-submodule of $A$. Let $K$ be the kernel of $G$ acting on $A/T$. Then $G_{0}=K\cap{\mathrm{CSp}}(2m,q)$ by Lemma 3.6. Since $G/G_{0}\leqslant{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)/{\mathrm{Sp}}(2m,q)\cong{\Gamma\mathrm{L}}(1,q)$, we have $K=G_{0}$. Hence, $G_{0}$ is the kernel of $G$ actign on $A/T$, as in part (ii).
Clearly that $S=T\oplus B$ is an ${\mathbb{F}}_{p}G$-submodule and $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/S=\dim A/T=f$. Hence $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$ if $S\subseteq W$. We assume that $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$. Note that each $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-submodule with dimension larger than $f$, which yields that $B\subseteq W$. If $T$ is not an ${\mathbb{F}}_{p}G$-submodule of $W$, then $T\cap W$ is an ${\mathbb{F}}_{p}G$-submodule of codimension no more than $f$. When $p\nmid m$, $T$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module by Lemma 3.6, which yields a contradiction. When $p\mid m$, $T\cap W\subseteq T_{0}$ given in Lemma 3.6 where $\dim T_{0}=f$, which is also a contradiction to $\dim T/(T\cap W)\leqslant f$. Thus, we obtain that $T\subseteq W$ and $S:=T\oplus B\subseteq W$, as in part (iii).
Assume that $G$ is transitive on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. If $G_{0}$ is not in the kernel of $G$ acting on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$, then $G$ induces a non-solvable permutation group acting on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. In this case, $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is an irreducible ${\mathbb{F}}_{p}G_{0}$-module with dimension equal to one of $n,2,4$ and $6$ by Theorem 3.1. We obtain that $m=1$, $2\mid f$ and $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\cong B_{f/2}$ as ${\mathbb{F}}_{p}G_{0}$-modules. In this case, the induced permutation group of $G_{0}$ on $B_{f/2}$ is isomorphic to $\Omega^{-}_{4}(p^{f/2})$ (see [18, page 45]), which is not transitive on non-zero vectors of $B_{f/2}$. Thus, $G_{0}$ acts trivially on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ and $B\subseteq W$. If $T$ is not an ${\mathbb{F}}_{p}G$-submodule of $W$, then $T/(T\cap W)$ is a trivial ${\mathbb{F}}_{p}G_{0}$-module, which is impossible by Lemma 3.6. We deduce that $S=B\oplus T\subseteq W$ and $G/G_{0}$ acting transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ since $G_{0}$ acting trivially on $A/T$. This also yields the converse direction, and part (iv) holds. ∎
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.

### v2

@@ 2.3. Extraspecial $q$-groups
Thus, $w^{g}\in W$ for any $w\in W$ and $g\in G$.
Hence, $G$ is a subgroup of ${\mathrm{Aut}}(N)^{V}$. By Lemmas 2.8 (i) and 2.11, the preimage $A$ of $G$ in ${\mathrm{Aut}}(N)$ is isomorphic to $p^{2mn^{2}}{:}{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$, and $A^{V}=G\cong{\mathrm{C}\Gamma\mathrm{Sp}}(2m,q)$.
@@ 3. Exterior Square of Transitive Linear Groups
Let $p$ be an odd prime, $V$ be a vector space over ${\mathbb{F}}_{p}$. To classify special $3$-orbit $p$-group, Lemma 2.8 (iii) and (iv) lead us to focus on the pairs $(G,W)$ such that $W<\Lambda_{{\mathbb{F}}_{p}}^{2}(V)$ is invariant under $G\leqslant{\mathrm{GL}}(V)$ and $G$ acts transitively on non-zero vectors of both $V$ and $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)/W$. It was shown in [25, Theorem 1.2] that if $p=2$ and $\dim V=\dim\Lambda_{{\mathbb{F}}_{p}}^{2}(V)/W$, then either $G$ is solvable or $G^{(\infty)}\cong{\mathrm{SL}}_{3}(2^{f})$ for some integer $f$. This result was applied to prove Gross’ conjecture on $2$-automorphic $2$-groups. For classifying $3$-orbit $p$-group of odd prime $p$, we will prove Theorem A in this section, which deals with the case where $p$ is odd and $\dim V$ has not to be equal to $\dim\Lambda_{{\mathbb{F}}_{p}}^{2}(V)/W$.
Let $V$ be an irreducible ${\mathrm{F}}G$-module for some $G\leqslant{\mathrm{GL}}(V)$ over finite field ${\mathrm{F}}$. Then ${\mathrm{E}}={\mathrm{End}}_{{\mathbb{F}}_{p}G}(V)$ is an extension field of ${\mathrm{F}}$. So $V$ can be regarded as an ${\mathrm{E}}G$-module, denoted by $U$ (see [1, Section 9]). Let $\rho:G\rightarrow{\mathrm{GL}}(U)$ be the homomorphism corresponds to the ${\mathrm{E}}G$-module $U$. For $\phi\in{\mathrm{Gal}}({\mathrm{E}}/{\mathrm{F}})$, denote by $U^{\phi}$ the ${\mathrm{E}}G$-module such that $u^{g}=(u^{\rho(g)})^{\phi}$ for $u\in U$ and $g\in G$. We have the following decomposition given in [25, Lemma 2.2]:
$\Lambda^{2}_{\mathrm{F}}(V)=A\oplus B\mbox{ with }B=\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}\mbox{, where}$ (1)
  1. (a)
    $A\cong\Lambda_{\mathrm{E}}^{2}(U)$ as ${\mathrm{F}}G$-modules;
  2. (b)
  3. (c)
We remark that a linear group $G\leqslant{\mathrm{GL}}(V)$ acts transitively on non-zero vectors is called a transitive linear group. The classification of transitive linear groups is a part of well-known classification of $2$-transitive groups, see [8, Table 7.3] for instance.
@@
Let $G\leqslant{\mathrm{GL}}(n,p)$ be a transitive linear group with odd prime $p$. Then one of the following statements holds.
@@
First, we consider the case where ${\mathrm{SL}}(m,q)\lhd G$. This case is similar to [25, Theorem 1.2], and we give a simple proof below for completeness.
@@
Let $V={\mathbb{F}}_{p}^{n}$, and let $G\leqslant{\mathrm{GL}}(V)$ such that $G^{(\infty)}\cong{\mathrm{SL}}(m,q)$ with $q=p^{f}$ odd, $n=mf$ and $m\geqslant 3$. Then $G$ acts transitively on non-zero vectors of $M=\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ if and only if $m=3$ and $M$ is isomorphic to the dual of $V$ as ${\mathbb{F}}_{p}G^{(\infty)}$-modules. In particular, $W$ is uniquely determined by $G^{(\infty)}$.
Proof.
Note that the induced linear group of $G^{(\infty)}$ on $M$ is either trivial or has a composition factor ${\mathrm{PSL}}(m,q)$. By (1), we can decompose $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ into a direct sum of ${\mathbb{F}}_{p}G^{(\infty)}$-submodules:
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B=A\oplus\left(\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}\right).$
Let $U$ be the natural ${\mathbb{F}}_{q}G^{(\infty)}$-module. Then $A$, ${\mathbb{F}}_{p}^{(\infty)}$-isomorphic to $\Lambda^{2}_{{\mathbb{F}}_{q}}(U)$, is an irreducible ${\mathbb{F}}_{p}G^{(\infty)}$-module, see [26, Theorem 2.2]. By Steinberg’s Twisted Tensor Product Theorem [33], each of $B_{j}$ is an irreducible ${\mathbb{F}}_{p}G^{(\infty)}$-module. Moreover, we have that $M$ cannot be a trivial ${\mathbb{F}}_{p}G^{(\infty)}$-module. Thus, the induced linear group of $G^{(\infty)}$ has a composition factor ${\mathrm{PSL}}(m,q)$. By Theorem 3.1, we have that $G^{(\infty)}$ acts faithfully and transitively on non-zero vectors of $M$ with $\dim M=\dim V=mf$. Then one of $A$ and $B_{j}$ for $j=1,...,\lfloor\frac{f}{2}\rfloor$ has dimension $n$. Remark that
$\dim A=f(\dim_{{\mathbb{F}}_{q}}\Lambda^{2}_{{\mathbb{F}}_{q}}(U))=\frac{m(m-1)}{2}f,$
$\dim B_{j}=f(\dim_{{\mathbb{F}}_{q}}U)^{2}=m^{2}f\mbox{ for $1\leqslant j<\frac{f}{2}$ and }$
$\dim B_{\frac{f}{2}}=\frac{f}{2}(\dim_{{\mathbb{F}}_{q}}U)^{2}=\frac{m^{2}f}{2}\mbox{ when $f$ is even}.$
Hence, we have that $M\cong A$ as ${\mathbb{F}}_{p}G^{(\infty)}$-modules and $m=3$. By [26, Theorem 2.2], $M$ is isomorphic to the dual of $V$ as ${\mathbb{F}}_{p}G^{(\infty)}$-modules.
Conversely, if $M$ is isomorphic to the dual of $U$ as ${\mathbb{F}}_{p}G^{(\infty)}$-modules when $m=3$. Then $\dim M=\dim V$, and hence the above arguments imply that $B=W$. It suffices to show that $B$ is an ${\mathbb{F}}_{p}G$-submodule. Since $G^{(\infty)}\lhd G$, $B^{g}$ is an ${\mathbb{F}}_{p}G_{(\infty)}$-submodule for $g\in G$. Suppose that $B^{g}\neq B$. Then $B^{g}/(B^{g}\cap B)\cong(B+B^{g})/B\cong A$ as ${\mathbb{F}}_{p}G^{(\infty)}$-modules. As $B^{g}$ is a completely reducible ${\mathbb{F}}_{p}G^{(\infty)}$-module, it contains an irreducible ${\mathbb{F}}_{p}G^{(\infty)}$-submodule $B_{0}$ such that $\dim B_{0}=\dim A=mf$. This is impossible since $B_{0}^{g^{-1}}$ is an irreducible ${\mathbb{F}}_{p}G$-submodule of $B$. Thus, we have that $B$ is an ${\mathbb{F}}_{p}G^{(\infty)}$-submodule for $g\in G$.
The “in particular” part follows since $W$ is always equal to $B$, which is determined by $G^{(\infty)}$. ∎
Now we deal with the case where ${\mathrm{Sp}}(2m,q)\lhd G$, which is the main difference between Theorem A and [25, Theorem 1.2].
Lemma 3.3.
Let $V={\mathbb{F}}_{p}^{n}$, and let $G\leqslant{\mathrm{GL}}(V)$ such that $G^{(\infty)}\cong{\mathrm{Sp}}(2m,q)$ with $q=p^{f}$ odd, $n=2mf$ and $m\geqslant 1$.
@@
    There exists a unique ${\mathbb{F}}_{p}G$-submodule $S$ of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ with codimension $f$ such that $G^{(\infty)}$ acts trivially on $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/S$.
@@
    If $G$ acts transitively on non-zero vectors of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$, then $W\supseteq S$.
@@
    If $\dim\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W\leqslant f$, then $W\supseteq S$.
@@
Let $U$ be the natural ${\mathbb{F}}_{q}G^{(\infty)}$-module. We decompose $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ into a direct sum of ${\mathbb{F}}_{p}G^{(\infty)}$-submodules as in (1):
@@
$\Lambda^{2}_{{\mathbb{F}}_{p}}(V)=A\oplus B=A\oplus\left(\bigoplus_{j=1}^{\lfloor\frac{f}{2}\rfloor}B_{j}\right).$
By Steinberg’s Twisted Tensor Product Theorem, we have that each of $B_{j}$ for $j=1,...,\lfloor\frac{f}{2}\rfloor$ is a faithful and irreducible ${\mathbb{F}}_{p}G^{(\infty)}$-module. Remark that $A\cong\Lambda_{{\mathbb{F}}_{q}}^{2}(U)$ has a ${\mathbb{F}}_{q}G^{(\infty)}$-submodule $T$ of codimension $1$ over ${\mathbb{F}}_{q}$ (see [10, 22, 28]) such that
  1. (a)
    if $p\nmid m$ then $T$ is an irreducible ${\mathbb{F}}_{p}G^{(\infty)}$-submodule;
  2. (b)
    if $p\mid m$ then $T$ is indecomposable such that $T/\mathrm{rad}(T)$ is an irreducible ${\mathbb{F}}_{p}G^{(\infty)}$-module, where $\dim_{{\mathbb{F}}_{q}}\mathrm{rad}(T)=1$.
Let $R$ be the radical of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ as ${\mathbb{F}}_{p}G$-module. Then either $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)$ is completely reducible when $p\nmid m$, or $R=T$ when $p\mid m$.
Part (i). It is clear that $S\supseteq B$, otherwise $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)/S$ is ${\mathbb{F}}_{p}G^{(\infty)}$-isomorphic to $B_{j}$ for some $j$. Thus, $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/S$ is naturally ${\mathbb{F}}_{p}G^{(\infty)}$-isomorphic to some quotient of $A$. Then we obtain that $S\supseteq T+B$. Note that $\dim A/T=f$, and thus $S=T+B$. It suffices to show that $S$ is an ${\mathbb{F}}_{p}G$-submodule. Suppose that $S^{g}\neq S$ for some $g\in G$. Then $S^{g}/(S\cap S^{g})\cong(S+S^{g})/S$ is a trivial ${\mathbb{F}}_{p}G^{(\infty)}$-module. This yields that $G^{(\infty)}$ acts trivially on $S/(S^{g^{-1}}\cap S)$. Remark that $\mathrm{rad}(S)=\mathrm{rad}(T)\oplus B$ contains no trivial ${\mathbb{F}}_{p}G^{(\infty)}$-submodules, a contradiction. Hence, $S$ is an ${\mathbb{F}}_{p}G$-module.
Part (ii). Suppose that $M$ is not a trivial ${\mathbb{F}}_{p}G^{(\infty)}$-module. Let $M=\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$. By Theorem 3.1, we have that $G^{(\infty)}$ acts faithfully and transitively on non-zero vectors of $M$ with $\dim M=\dim V$. Since $M$ is ${\mathbb{F}}_{p}G^{(\infty)}$-isomorphic to a submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/R$, it yields that $M$ is ${\mathbb{F}}_{p}G^{(\infty)}$-isomorphic to either $B_{j}$ for some $j$ or $T$ when $p\nmid m$. Note that
$\dim T=f\left(\frac{2m(2m-1)}{2}-1\right),$
$\dim B_{j}=f(\dim_{{\mathbb{F}}_{q}}U)^{2}=4m^{2}f\mbox{ for $1\leqslant j<\frac{f}{2}$ and }$
$\dim B_{\frac{f}{2}}=\frac{f}{2}(\dim_{{\mathbb{F}}_{q}}U)^{2}=2m^{2}f\mbox{ when $f$ is even}.$
By comparing dimensions, we have that $M$ is ${\mathbb{F}}_{p}G^{(\infty)}$-isomorphic to $B_{\frac{f}{2}}$ with $m=1$ and $f$ is even. Recall that $2\cdot B_{\frac{f}{2}}\cong U\otimes U^{\phi^{\frac{f}{2}}}$ and $G^{(\infty)}\cong{\mathrm{SL}}(2,q)$. Then the induced linear group of $G^{(\infty)}$ acting on $2\cdot B_{\frac{f}{2}}$ is isomorphic to $\Omega^{-}_{4}(p^{f/2})$ (see. [21, page 45]), which is not transitive on non-zero vectors. Thus, we have that $M$ is a trivial ${\mathbb{F}}_{p}G^{(\infty)}$-module.
Note that $B$ is a completely reducible ${\mathbb{F}}_{p}G^{(\infty)}$ with no trivial ${\mathbb{F}}_{p}^{(\infty)}$-submodules. This implies that $W\supset B$. Thus, we have that $S=T\oplus B\subseteq W$ when $p\mid m$ since $T=R$ is the radical of $\Lambda_{{\mathbb{F}}_{p}}^{2}(V)$. When $p\nmid m$, we have that $T$ is a non-trivial irreducible ${\mathbb{F}}_{p}^{(\infty)}$-submodule of $A$. Thus, $S=T\oplus B\subseteq W$ as desired.
Part (iii). Note that each non-trivial ${\mathbb{F}}_{p}G^{(\infty)}$-submodule of $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/R$ has dimension larger than $f$. Then $\Lambda^{2}_{{\mathbb{F}}_{p}}(V)/W$ is trivial. Therefore, $S\subseteq W$ since $S/(S\cap R)$ has no trivial submodules. ∎
Now we are ready to give a proof for Theorem A.
Proof of Theorem A.